OpenStudy (anonymous):

the cube root of 108x^5 ------ = the cube root of 4x^2

OpenStudy (lgbasallote):

do you mean \[\huge \frac{\sqrt[3]{108x^5}}{\sqrt[3]{4x^2}} = ?\]

OpenStudy (anonymous):

yes

OpenStudy (perl):

this program is buggy,

OpenStudy (perl):